### v2

@@ 4 Application 1: Motion transfer
Our first application is motion transfer. Motion transfer allows creators to control the motion in a fine way, which cannot be achieved using text. In this problem we are given a video $V_{S}$ that contain some pattern of movement, and we wish to generate a new video $V_{T}$ that follows the same movement patterns. The movement can be of entities in the video, or the camera or both.
Figure 4: (click-to-view-online) Qualitative Results, Motion Transfer. Each pair of rows are frames from source (pair-top) real video and target (pair-bottom) generated video. Transferring Motion by Queries allows to use source videos to inject camera motion (top), non-rigid movement (middle), and combinations of movements (bottom). See the supplemental for videos and more examples.
@@
4.1 Method for Motion transfer
Our approach, called “Motion by Queries” follows the experiment described in section 2. First, we extract a series of Q-features from $V_{S}$, by denoising the model to various timesteps ($t=1000,980,\ldots,600$), obtaining $[Q_{S}(50),\ldots,Q_{S}(30)]$. Then, we inject those Q features during the generation of the target video $V_{T}$, this time with the new prompt $\tau_{T}$.
Implementation details: We only record and inject the Q features of the conditional forward pass. To ensure consistency with the video model’s latent space, we follow Meiri et al. (2023) and add minor noise to the source video latent (two steps) and denoise it back to a new clean latent before extracting Q-features. For recording Q features, we use the same noise seed across different noise levels. During generation, we start from a random initialization. For our experiments, we use VideoCrafter2 (Chen et al., 2024).
4.2 Experiments
We evaluate Motion by Queries qualitatively and quantitatively using a Motion-Transfer benchmark Wang et al. (2024a), comprising 66 video-edit text pairs, guided by 22 source videos. Videos were sourced from DAVIS Perazzi et al. (2016), WebVID Bain et al. (2021), and online resources, representing diverse scenes, objects, and motion. For the base text-to-video model, we employ VideoCrafter2 Chen et al. (2024).
Baselines: We compare our approach with recent methods: (1) Diffusion Motion Transfer (DMT), a test-time optimization approach Yatim et al. (2024), and (2) Motion-Inversion (MI) Wang et al. (2024a), a fine-tuning-based concurrent method that learns motion-specific embeddings that modulate temporal attention activations from the source video’s motion. MI indicates results we reproduced from their online code. Except for one case in the supplemental, all qualitative comparisons use videos from MI and DMT project pages. We also adopt from Wang et al. (2024a) the quantitative results of Video Motion Customization (VMC) Jeong et al. (2024) and Motion-Director (MD) Zhao et al. (2024), both are fine-tuning-based approaches.
Evaluation Metrics: We measure Identity Leak (ID. Leak) from source to target video, as described in Section 3. Following Wang et al. (2024a), we also measure Motion Fidelity (MF), Text Similarity (Text), and Temporal Consistency (Temp. C.). MF measures point track correlation between generated and source videos, as described in Section 3. Text Similarity is the average CLIP-Text score between generated frames and the textual prompt Radford et al. (2021). Temporal Consistency Huang et al. (2024) measures video coherence by averaging CLIP-Image similarity between consecutive frames in the generated video.
@@ 4.2 Experiments
Additionally, we compare the runtimes of different methods and their relative overhead compared to the base video model each method uses. We break down the runtime into its components: “Source Invers.” is the time for DDIM inversion in baselines or recording Q features in our approach. “Optim./Tuning” is the optimization or fine-tuning time, and “Infer” is the time to generate a target video. Note that in DMT, Infer and Optim are done jointly. “Sum” is the total time for each method, and “Overhead” is the ratio between total time and inference time in the base model. Finally, runtimes were evaluated using an NVIDIA H100 GPU, when generating a video of 24 frames at a resolution of 576$\times$320.
@@
Quantitative Results: In terms of performance, our approach demonstrates competitive results across various metrics while maintaining simplicity and optimization-free advantages (Tables 1, 2). Specifically, our method achieves lower (better) identity leakage (38.6) compared to MI (43.7), while text alignment (28.8) and temporal consistency (97.0) perform on par with baselines. Regarding motion fidelity (MF), our method achieves 91.5. Although lower than MI (97.0), fidelity remains adequate for many practical applications, as supported by the qualitative results. Notably, our method excels in computational efficiency. Compared to MI and DMT requiring 208 and 410 seconds with $\times$23-45 overhead, our method runs in 70 seconds with minimal $\times$1.2 overhead to the base model (VC2).
@@
Id. Leak $\downarrow$ Text $\uparrow$ MF $\uparrow$ Temp. C. $\uparrow$
DMT - 28.8 78.8 93.6
VMC - 27.1 93.7 94.6
MD - 30.4 93.9 93.3
MI - 31.1 95.5 93.5
MI 43.7 $\pm$ 1.5 29.2 $\pm$ 0.6 97.0 $\pm$ 0.4 96.8 $\pm$ 0.2
Ours 38.6 $\pm$ 1.8 28.8 $\pm$ 0.6 91.5 $\pm$ 0.9 97.0 $\pm$ 0.2
@@
Table 1: Quantitative Evaluation Metrics, Motion Transfer. Values are reported as mean $\pm$ standard error of the mean (S.E.M).
Time Source Optim./ Infer Sum Over-
[sec.]: Invers. Tuning head $\downarrow$
@@
Table 2: Runtime Comparison. Existing methods are $\times$23-45 slower than their base model. Ours adds merely $\times$1.2 overhead.
Qualitative Results: In Figure 4, we show that Motion by Queries helps guide generated videos to match the motion in source videos, including camera motion (top row), non-rigid object movement (middle row), and combinations of both types of motion (bottom row). While for quantitative results, we used Q injection ending at $t=600$. For qualitative results, we selected videos with Q injection ending between $t=800$ and $t=600$: closer to $t=800$ for camera motion and closer to $t=600$ for non-rigid movement. In Figure 5, we compare our results to MI and DMT. In the monkeys example, our method generates motion that resembles the source video, though it is with slightly lower motion magnitude compared to the baselines. For the boy and train examples, our results are comparable to those of MI and DMT. Additional comparisons reveal cases where the MI baseline suffers from identity leakage, aligning with the quantitative results. The supplemental also includes further comparisons and more examples of custom camera motion, and hybrid combinations of object and camera motion.
@@ 5 Application 2: Consistent multi-shot video generation
Generating videos from text prompts has progressed significantly, but current methods still struggle with maintaining long, coherent video sequences. A promising approach is to generate multiple short videos featuring the same characters, as cinematic videos often consist of shorter shots. The challenge lies in ensuring character consistency across scenes, as current T2V models excel at generating individual clips but struggle with consistency across multiple shots. Our goal is to generate multiple video shots with consistent characters, preserving their identity and motion dynamics across all scenes. This would open up new possibilities in content creation, storytelling, and entertainment.
Previous work  (Tewel et al., 2024; Fan et al., 2024) has focused on consistent character generation in the text-to-image domain by extending the self-attention mechanism, allowing images to share visual features. However, as shown in our analysis, video features often encode both identity and motion, leading to challenges when applying these methods to video. Naive extended attention mechanisms can cause synchronized or diminished motion, which makes image-based consistency methods unsuitable for video.
Building on this analysis, we propose a solution tailored for consistent video generation  [^1] The methods and results in this section are based on the arXiv version 1 (v1) Atzmon et al. (2024) of this work. Here, in version 2 (v2), we extend and further analyze those findings.. Our method preserves character identity and motion dynamics by combining extended attention for consistency with a vanilla, non-consistent text-to-video step to maintain dynamic motion. In the following, we describe the key modifications we introduced to Tewel et al. (2024) for consistent video generation.
ConsiStory Tewel et al. (2024) (See extended overview in the appendix, Sec. B.2) applies extended attention by using a subject mask, allowing images to share K, V self-attention features to maintain subject consistency. This reduces layout variability in the generated images. Then, to restore layout diversity, ConsiStory injects Q features from a vanilla, non-consistent sampling step. However, while this method improves diversity in images, it does not work well for videos, where the injected Q features fail to maintain the desired consistency and diversity across frames.
@@
The naive extension of the extended attention approach to videos involves an extended self-attention mechanism that allows frames from different videos to attend to each other. However, this leads to reduced motion within each video. We hypothesize that self attention primarily ensures coherence within a video-shot, so extending it to multiple video-shots causes the different video-shots to become overall aligned, reducing the motion. Since the video Q-features are strongly tied to layout and motion, one approach to preserve each video’s original motion (i.e., pre-extended attention) is to inject Q-features from its vanilla version. While this helps preserve motion diversity, it also results in identity leakage from the original video, breaking our goal of maintaining character consistency across videos, as shown in Fig. 3.
To address this, we propose a two-stage solution for motion injection. Q preservation followed by Q Flow. In the early diffusion steps, which primarily control motion and layout, we inject the Q-features from vanilla videos generated without extended attention.
Then in the subsequent Q-Flow phase we apply a relaxed version of our Q-injection that preserves the flow of Q-features rather than injecting the features directly. This technique is inspired by TokenFlow Geyer et al. (2023), originally designed for T2I models.
Specifically, for each vanilla video, we first compute the nearest-neighbor correspondence field on its Q-features, defining a Q-feature flow that we aim to maintain in the generated consistent video. We then inject this correspondence into our generated video. For the exact definition of our Q Flow injection, see Appendix B.12.
@@ 5.1 Experiments
To demonstrate the importance of the steps suggested above, we conducted a study to investigate the effects of self-attention query (Q) tokens on motion and identity for consistent video generation. In the Appendix (Sec. B), we further show extensive evaluation including comparisons to baselines, a user study, and an ablation study. Fig. 6 illustrates typical generations for different interventions on Q tokens when combined with the extended self-attention mechanism of Fig. B.3. When we do not intervene in the Q tokens (Fig. 6, 4th row - “No Q Intervention”), subject identity is well-maintained across video shots, but motion quality significantly degrades. This manifests in: 1) Motion synchronization: movements become synchronized across video shots, e.g., the dog’s head turning simultaneously in all shots. 2) Reduced variability in motion style and pose: similar actions are repeated across shots, e.g., the dog’s leap, the Muppet’s centered swaying, the camera movement becomes static in the skating Muppet shot. 3) Motion artifacts: to reconcile the reduced motion variability with each scene’s text prompt, videos tend towards motion-artifacts. For example, the skating Muppet’s body appears frozen while its legs are displaced to visually accommodate the “skating” action. In contrast, combining extended attention with injected Q tokens cached from vanilla diffusion-sampled video shots (Fig. 6, 3rd row - “Full Q Preservation”) restores motion but largely loses subject identity. For instance, the Muppet’s colors revert to those of the vanilla model.
These observations reiterate our main finding about the dual nature of Q tokens. Injecting vanilla Q tokens restores motion, showing their influence on movement. Simultaneously, it leads to a leakage of the vanilla identity into the multi-shot consistency mechanism. This causes a loss of the shared subject identity (e.g., the Muppet’s color change).
Our approach (Fig. 6, 1st row - “Ours”) achieves a balance between both worlds. By intervening in the Q tokens throughout generation, we restore most of the original motion, including nuanced details like body and face orientations, postures, and natural movement of specific body parts (dog’s ears, Muppet’s hands and legs). Even the parallax style of video shots is preserved (right Muppet video shot). Our method’s effectiveness stems from our two-step process. First, Q preservation in early denoising steps establishes motion structure before identity is fully set. Then, flow-based Q injection allows Q values to evolve to better match the novel (more consistent) generation while enforcing some structural alignment through the flow process.
6 Limitations
In motion transfer, our method achieves slightly average motion fidelity compared to existing approaches, particularly with rapid movements (Figure 7 Appendix, top row). When transferring motion between different object classes, the source object’s shape can occasionally influence the target, such as when a pirate ship takes on characteristics of the source swan, or when a Colosseum inherits spatial properties from a mug (Figure 7).
In multi-shot consistent generation, balancing identity preservation and motion quality is still challenging. We find that Q injection may be too strong and still hurt identity. To manage this, motion preservation can be compromised by partially dropping out Q injection (See Section B.14).
7 Conclusion
Our analysis reveals key differences in how Q-injection behaves in text-to-video versus text-to-image models. Videos require more denoising steps to establish motion patterns, and Q vectors encode both motion and identity information, extending beyond their traditionally understood role in image models. When using extended attention for multi-shot video generation, motion tends to synchronize across videos, requiring longer Q-injection periods that affect both motion and identity. These insights enabled two practical applications: an efficient zero-shot motion transfer method and a training-free approach for consistent multi-shot videos. In a broader perspective, these findings challenge the separation of ”where” and ”what” pathways established in text-to-image models Tewel et al. (2023); Patashnik et al. (2023); Alaluf et al. (2023), providing novel evidence that queries in video models influence both motion (structure over time) and identity, suggesting a more complex interplay between these pathways than previously appreciated.
7.1 Broader Impact
This work advances text-to-video generation, offering applications in creative and entertainment industries. However, it introduces ethical challenges, including the potential misuse for creating misleading or harmful content. Future societal impacts could involve both enhanced accessibility to video creation and increased risks of misinformation.
@@ Appendix A Appendix
Figure 7: Limitations. Cases where the source subject shape affects the target object.
A.1 Background: Self-Attention in T2V models

### v3

@@ 4 Application 1: Motion transfer
Our first application is motion transfer. It allows creators to control motion in a fine way, which cannot be achieved using text. In this problem we are given a video $V_{S}$ that contain some pattern of movement, and we wish to generate a new video $V_{T}$ that follows the same movement patterns. The movement can be of entities in the video, or the camera or both.
Our approach, called “Motion by Queries” follows the experiment described in section 2. First, we extract a series of Q-features from $V_{S}$, by denoising the model to various timesteps ($t=1000,980,\ldots,600$), obtaining $[Q_{S}(50),\ldots,Q_{S}(30)]$. Then, we inject those Q features during the generation of the target video $V_{T}$, this time with the new prompt $\tau_{T}$. For WAN 2.1, we extracted Q-features from a single low-noise timestep and injected them into all higher-noise steps during target generation, inspired by Ling et al. (2025). See appendix for additional implementation details.
4.1 Experiments
We evaluate Motion by Queries qualitatively and quantitatively using a Motion-Transfer benchmark Wang et al. (2024a), comprising 66 video prompts, guided by 22 source videos. Videos were sourced from DAVIS Perazzi et al. (2016), WebVID Bain et al. (2021), and online resources, representing diverse scenes, objects, and motion. For the base text-to-video model, we use VideoCrafter2 (VC2).
Baselines: We compare our approach with recent methods: (1) Diffusion Motion Transfer (DMT), a test-time optimization approach Yatim et al. (2024), a (2) Motion-Inversion (MI) Wang et al. (2024a), a fine-tuning method that learns motion-specific temporal embeddings from a source video, and (3) MotionClone (MC) Ling et al. (2025), a zero-shot approach where motion is guided by temporal transformer self-attention maps extracted when U-Net processes source videos.  MI, MC indicate results we reproduced from their online code. Except for one case in the supplemental, all qualitative comparisons use videos from MI and DMT project pages. We also adopt from Wang et al. (2024a) the quantitative results of Video Motion Customization (VMC) Jeong et al. (2024) and Motion-Director (MD) Zhao et al. (2024), both are fine-tuning-based approaches.
Evaluation Metrics: We measure Motion Fidelity (M. Fidel.), the point track correlation between generated and source videos, as described in Sec. 3; Temporal Flicker (T. Flick.) which quantifies flickering artifacts, using mean absolute error between consecutive frames, as in VBench Huang et al. (2024); Text Similarity (Text), the average CLIP-Text score between generated frames and the textual prompt Radford et al. (2021); Temporal Consistency (Temp. C.) which measures video coherence by averaging CLIP-Image similarity between consecutive frames  Huang et al. (2024); Identity Leak (Id. Leak) to quantify the similarity between source and target video frames, as described in Sec. 3; Finally, Aesth., Smooth., and Bk Cons. which are VBench quality metrics for visual appeal, temporal smoothness, and background consistency.
@@ 4.1 Experiments
Additionally, we compare runtimes and overhead relative to each method’s base model. We break the runtime to components: “Invers.” is for inversion or feature recording from source video, “Optim.” for optimization or tuning, “Infer” for sampling a new video, “Sum” is the total runtime, and “Overhead” is the ratio of “Sum” to inference in the base model. Runtimes were measured on NVIDIA H100, when generating at a resolution of 576$\times$320.
@@
Quantitative Results: Our approach demonstrates competitive results across various metrics while remaining simple and optimization-free (Tables 1, 2). Specifically, our method achieves lower (better) identity leakage (38.6) than MI (43.7), with text alignment (28.8) and temporal consistency (97.0) comparable to baselines. For motion fidelity (MF), our method reaches 91.5, which, while lower than MI (97.0), fidelity remains sufficient for some practical scenarios as supported by qualitative results. Compared to MC, our method attains higher identity leakage (38.6 vs. 24.2, lower is better), while achieving improved temporal stability, as reflected in a higher flickering score (95.1 vs. 86.0).  In terms of efficiency, our method requires only $\times$1.2 overhead (70s) relative to the base model (VC2), and is more efficient than MC ($\times$12, 104s), MI ($\times$23, 208s), and DMT ($\times$45, 410s).
@@
M. Fidel. $\uparrow$ T. Flick. $\uparrow$ Text $\uparrow$ Temp. C. $\uparrow$ Id. Leak $\downarrow$ Aesth. $\uparrow$ Smooth. $\uparrow$ Bk Cons. $\uparrow$
DMT 78.8 - 28.8 93.6 - - - -
VMC 93.7 - 27.1 94.6 - - - -
MD 93.9 - 30.4 93.3 - - - -
MI 95.5 - 31.1 93.5 - - - -
MI 97.0 $\pm$ 0.4 92.2 $\pm$ 0.4 29.2 $\pm$ 0.6 96.8 $\pm$ 0.2 43.7 $\pm$ 1.5 52.6 $\pm$ 1.2 95.6 $\pm$ 0.3 94.5 $\pm$ 0.3
MC 95.0 $\pm$ 0.7 86.0 $\pm$ 0.6 29.9 $\pm$ 0.4 95.8 $\pm$ 0.3 24.2 $\pm$ 1.1 55.6 $\pm$ 0.9 93.5 $\pm$ 0.4 93.3 $\pm$ 0.4
Ours 91.6 $\pm$ 0.9 95.1 $\pm$ 0.3 28.8 $\pm$ 0.6 97.0 $\pm$ 0.2 38.6 $\pm$ 1.8 54.1 $\pm$ 1.3 97.3 $\pm$ 0.2 94.9 $\pm$ 0.4
@@
Table 1: Quantitative Evaluation Metrics, Motion Transfer. Values are mean $\pm$ S.E.M.
Time [sec.] Invers. Optim. Infer Sum Overhead $\downarrow$
@@
Table 2: Runtime Comparison. Existing methods are $\times$12-45 slower than their base model. Ours adds merely $\times$1.2 overhead.
Qualitative Results: In Figure 4 and supplemental, we show that Motion by Queries guides videos generated by VC2 to follow the motion in source videos, capturing both camera motion (top), and non-rigid object movement (bottom). Figure 5 compares our results with MC, MI, and DMT: in the monkeys example, our method produces motion that resembles the source video, albeit with slightly lower motion magnitude than the baselines. For the boy and train examples, our results are comparable to those of MI and DMT. Supplemental comparisons further reveal cases where MC exhibits temporal instability  and MI identity leakage, as seen in the quantitative results. They also provide examples illustrating custom camera motion and hybrid object-camera motion. Figure 6 and supplemental present results with the WAN 2.1 DiT model. Compared to VC2, WAN exhibits a stronger physical prior: it tends to generate motion that is physically plausible and realistic, sometimes prioritizing physical accuracy over exact pixel-level replication. This is noticeable in non-rigid motion scenarios, such as the wall climbing example. We also find that WAN is more sensitive to the choice of initial noise, resulting in greater variability in the generated outputs.
Figure 6: (click-to-view-online) Qualitative Results with Wan 2.1, Motion Transfer. Each row pair shows source (top) and generated (bottom) frames. See supplemental for videos and more examples.
@@ 5 Application 2: Consistent multi-shot video generation
Generating long, coherent sequences remains difficult for text-to-video models, yet this ability is key for rich storytelling and creative video applications. A practical approach is to generate several short video shots with the same characters, as in cinematic videos. The main challenge is keeping character consistency across shots, since current models excel at single clips but struggle with cross-shot generation.
Prior work (Tewel et al., 2024; Fan et al., 2024) tackled character consistency in text-to-image by sharing self-attention features. However, as our analysis shows, video features encode both identity and motion, making it challenging to apply these methods to video, where naive extended attention can cause synchronized or diminished motion. To address this, we propose a method tailored for consistent video generation, combining extended attention for consistency Tewel et al. (2024) with a two-step Q injection process. Early Q preservation sets motion structure, followed by flow-based Q injection that lets Q values evolve for better consistency while maintaining structure.
@@
We build on ConsiStory Tewel et al. (2024) (Appendix  B.2) that applies extended attention within a subject mask to share K, V self-attention features, but reduces layout variability. To restore diversity, ConsiStory injects Q features from a vanilla, non-consistent sampling step. While this works for images, injecting Q features from vanilla videos preserves motion diversity but causes identity leakage from those videos, breaking character consistency (Fig. 3).
Our two-stage solution—Q preservation followed by Q Flow—addresses this. In early diffusion steps, which primarily control motion and layout, we inject the Q-features from vanilla videos generated without extended attention. Then, in the subsequent Q-Flow phase we apply a relaxed Q-injection that preserves the flow of Q-features rather than injecting the features directly. This technique is inspired by TokenFlow Geyer et al. (2023). Specifically, for each vanilla video, we first compute the nearest-neighbor correspondence field on its Q-features, defining a Q-feature flow to maintain in the generated consistent video. We then inject this correspondence into our generated video (see Appendix B.11 for details).
@@ 5.1 Experiments
We investigated the effects of self-attention query (Q) tokens on motion and identity for consistent video generation described above. Further extensive evaluations, including baselines, user studies, and ablations, are in Appendix B. Fig. 7 shows typical generations for different Q token strategies when combined with extended self-attention.
Without Q intervention (Fig. 7, 4th row), subject identity is preserved across shots, but motion quality drops: (1) movements become synced, e.g., the dog’s head turning simultaneously in all shots. (2) motion style and pose vary less: repeatition across shots, e.g., the dog’s leap, the Muppet’s centered swaying, the camera movement becomes static in the skating Muppet shot. (3) motion-artifacts, to reconcile motion sharing with the varying text prompts, e.g., the skating Muppet’s body appears frozen while its legs are displaced. In contrast, Injecting Q tokens from vanilla videos (3rd row) restores motion but loses subject identity, e.g. the Muppet’s colors revert to those of the vanilla model. This reiterate the dual role of Q tokens: injecting vanilla Qs restores motion, but leaks the vanilla non-consistent identities into the multi-shot mechanism. Our approach (1st row) balances both, restoring the original motion, including nuanced details like body and face orientations, postures, specific body parts , as well as parallax camera movement.
6 Conclusion and Limitations
Our findings show that, unlike in text-to-image models, Q-injection in video models affects both motion and identity, requiring more denoising steps to establish motion patterns. This challenges the classic separation of ”where” and ”what” pathways Tewel et al. (2023); Patashnik et al. (2023); Alaluf et al. (2023), revealing a more complex interplay in video models. These insights led to two applications: zero-shot motion transfer and training-free consistent multi-shot video generation. Our method has limits: motion fidelity can drop with fast movements, and cross-class transfer may cause shape leakage (See Fig. 9 in Appendix). In multi-shot generation, balancing identity and motion remains hard—Q injection can still hurt identity, and partial dropout may be needed to preserve motion (Section B.13).
@@ Appendix A Appendix
Figure 8: (click-to-view-online) Identity Leakage and Motion Transfer with Other Text-to-Video Models.
Figure 9: Limitations. The source subject shape may affects the target object.
A.1 Motion Transfer Implementation Details
  • We record and inject only the Q features from the conditional forward pass.
  • When recording Q features, we use the same noise seed across different noise levels. During generation, however, we start from a random initialization.
  • In VideoCrafter2, to ensure consistency with the video model’s latent space, we follow Meiri et al. (2023): we add minor noise to the source video latent (two steps) and then denoise it back to a clean latent before extracting Q features.
  • For quantitative results, we used Q injection ending at $t=600$. For qualitative results with VideoCrafter2, we selected videos with Q injection ending between $t=800$ and $t=600$: closer to $t=800$ for camera motion, and closer to $t=600$ for non-rigid movement.
A.1.1 DiT-Based Models
  • For both WAN and LTX-Video, we used a 50-step flow-matching denoising scheduler as described by Esser et al. Esser et al. (2024). This schedule shifts the timestep allocation so that more steps are concentrated in the high-noise region. Specifically, we used the FlowMatchEulerDiscreteScheduler from huggingface, with their default $\mu=3.065$ hyperparam.
  • Similar to VideoCrafter2, to transfer the full magnitude of motion we had to inject Q features for a substantial amount of steps. For WAN, we injected Q features for $58\%$ or $60\%$ of the denoising schedule; for LTX-Video, we used $40\%$.
  • In WAN, we inject Q features only in layers 20–30. In all other models, Q injection is applied to all layers.
  • Current motion transfer benchmarks consist of short 16- or 24-frame videos, which is significantly shorter than the standard length of WAN 2.1 videos (81 frames). Therefore, for motion transfer in WAN, we repeat each frame twice and pad the last frame one additional time, therefore mapping $16\rightarrow 33$ and $24\rightarrow 49$. Accordingly, we double the frame rate of the generated videos to keep the duration the same as in the source video.
  • For LTX-Video only, we found that injecting Q features between different initial noise seeds preserved the findings about identity leakage but introduced visual artifacts. To mitigate this, we used identical seeds for different prompts and matched the global statistical moments of the Value features, improving compatibility with the injected Query features.
A.2 Background: Self-Attention in T2V models